A biochemist has 100 mL of a 0.10 M solution of a weak acid with pKa of 6.3. She adds 6.0 mL of 1.0 M HCL which changes the pH to 5.7. What was the pH of the original solution?
1
Expert's answer
2017-10-09T14:49:06-0400
Find hydrogen ion concentration of the original solution using the law of mass conservation after some further algebraic manipulation:
[H+]2+Ka[H+]-KaCa=0 (Ka - weak acid dissociation constant, Ca – weak acid concentration)
Ka=10-6.3=5.01*10-7
Set up the quadratic equation: [H+]2+5.01*10-7*[H+]-5.01*10-7 *0,10=0
Solve the quadratic equation: [H+]=2*10-4
pH=3.7